OpenStudy (anonymous):

suppose f(x)= x^4 +ax^2. What is the value of a if f has a local min at x=2?

OpenStudy (anonymous):

suppose . [f(x)=x^{4}+ax ^{2.}\] What is the value of a if f has a local min at x=2?

OpenStudy (anonymous):

im not sure if im supposed to find the derivative then set x = 2 then solve for a or something else..

OpenStudy (anonymous):

If there is a local min at x=2, then f'(2)=0.

OpenStudy (anonymous):

If it has a local min, then we know the derivative at that point must be 0. So take the derivative, set x = 2, and set the derivative =0, then solve for a.

OpenStudy (anonymous):

Which is essentially what Xavier said, but with more words ;p

OpenStudy (anonymous):

okay, i did that.. the answer is -8 but i got -4

OpenStudy (anonymous):

Did you get 4x^3+2ax=0 as the derivative?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Substitute x=2 4(2)^2+4a=0

OpenStudy (anonymous):

4(2)^3+4a=0 I mean

OpenStudy (anonymous):

it would be 4(2)^3

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

4a=-32

OpenStudy (anonymous):

ohhh. okay. wow.

OpenStudy (anonymous):

thank soo much. im stupid..

OpenStudy (anonymous):

Everyone has those moments :P

OpenStudy (amistre64):

tag!!

OpenStudy (amistre64):

tag